OpenStudy (anonymous):

What is the probability of randomly selecting a heart from a standard deck of 52 cards?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

13 hearts, 13 spades, 13 clubs, 13 diamonds = 52 total 13/52 = 1/4 = 0.25 probability

OpenStudy (anonymous):

thank you